# From Embeddings to Accuracy: Comparing Foundation Models for Radiographic Classification

### v1

@@ 3 Results
KNN F 91.2 90.1 94.5 92.6 96.2 90.1 96.4 93.0
M 85.2 88.8 90.2 93.5 95.3 87.4 96.6 91.0
Difference 6.0 1.3 4.3 -0.8 0.9 2.7 -0.2 2.0
LG F 90.4 91.0 95.7 92.7 98.0 91.5 97.7 93.8
M 89.8 89.6 93.5 95.1 96.2 90.6 97.6 93.2
Difference 0.5 1.4 2.2 -2.4 1.8 0.9 0.1 0.7
SVM F 92.2 91.5 94.6 94.9 98.0 91.6 98.0 94.4
M 88.5 91.1 94.5 95.5 96.2 90.6 98.1 93.5
Difference 3.7 0.4 0.1 -0.6 1.8 1.0 -0.1 0.9
RF F 87.9 91.2 95.1 92.9 97.5 91.2 97.4 93.3
M 86.5 90.0 93.4 95.0 94.7 89.7 97.0 92.3
Difference 1.5 1.2 1.8 -2.1 2.8 1.5 0.4 1.0
MLP F 91.7 91.1 95.4 91.3 97.7 91.7 97.9 93.8
M 90.6 90.8 93.0 94.5 96.7 91.0 98.1 93.5
Difference 1.1 0.3 2.3 -3.2 1.0 0.7 -0.3 0.3
    KNN = k-nearest neighbors, LG = logistic regression, SVM = support vector machine, RF = random forest, MLP = multi-layer perceptron, ETTube=endotracheal tube, NGTube=nasogastric tube, RFrac=rib fracture, VLine=vascular lines, F=female patients (709 in total), M=male patients (1059 in total).
Table 3: Age fairness assessment of AUC performance (%) from adapters trained by MedImageInsight embeddings
@@
KNN (0, 20] 98.7 84.3 100.0 99.1 100.0 79.3 99.7 94.4
(20, 40] 85.8 89.1 77.3 90.9 96.8 88.2 97.1 89.3
(40, 60] 86.7 91.0 92.7 94.9 95.5 90.0 97.3 92.6
(60, 80] 86.7 87.6 92.2 91.9 96.2 88.9 93.7 91.0
(80, 100] 92.9 86.2 95.5 86.3 96.3 85.4 98.1 91.5
std 5.5 2.6 8.5 4.8 1.8 4.3 2.2 1.9
LG (0, 20] 100.0 87.2 100.0 99.3 99.7 84.3 100.0 95.8
(20, 40] 91.1 91.1 93.9 91.9 98.5 91.4 97.9 93.7
(40, 60] 87.4 90.8 94.7 96.0 97.3 91.4 98.2 93.7
(60, 80] 90.4 88.4 92.8 95.1 97.3 90.7 96.0 93.0
(80, 100] 94.3 87.3 97.0 53.5 96.3 92.0 96.3 88.1
std 4.8 1.9 2.9 19.0 1.3 3.2 1.6 2.9
SVM (0, 20] 100.0 85.6 100.0 99.4 99.9 82.8 100.0 95.4
(20, 40] 87.5 90.4 93.7 94.7 97.9 90.5 98.2 93.3
(40, 60] 87.3 91.4 94.9 94.8 97.5 91.4 98.8 93.7
(60, 80] 91.1 90.9 93.4 95.1 97.4 91.6 96.5 93.7
(80, 100] 97.0 86.8 96.4 84.7 96.8 91.8 97.3 93.0
std 5.7 2.6 2.7 5.4 1.2 3.8 1.3 0.9
RF (0, 20] 99.3 87.8 100.0 99.3 100.0 85.0 100.0 95.9
(20, 40] 87.7 89.9 93.0 95.1 97.2 90.5 97.3 93.0
(40, 60] 83.7 91.3 94.6 95.2 96.7 90.5 98.2 92.9
(60, 80] 86.8 89.1 92.0 92.2 97.3 91.1 95.1 92.0
(80, 100] 92.5 87.9 99.5 80.0 94.2 93.4 96.8 92.0
std 6.1 1.5 3.7 7.3 2.1 3.1 1.8 1.6
MLP (0, 20] 100.0 86.9 100.0 99.6 99.7 84.6 100.0 95.8
(20, 40] 91.0 92.1 93.7 90.0 98.9 91.9 98.2 93.7
(40, 60] 87.8 91.7 93.5 95.4 97.2 91.0 98.6 93.6
(60, 80] 92.1 89.6 92.7 94.5 97.3 91.8 96.8 93.5
(80, 100] 94.9 87.4 96.1 58.0 95.9 91.3 97.0 88.6
std 4.6 2.4 3.0 16.8 1.5 3.1 1.3 2.7
    ETTube = endotracheal tube, NGTube = nasogastric tube, RFrac = rib fracture, VLine = vascular lines, std = standard deviation. A total of 154 patients with age in the range of (0, 20], 237 patients with age in the range of (20, 40], 570 patients with age in the range of (40, 60], 610 patients with age in the range of (60, 80], and 99 patients with age in the range of (80, 100] are involved in the analysis. Patients older than 100 are unidentified and are excluded from analysis.
Beyond fairness, computational efficiency was assessed through the training and inference times of each adapter model on a CPU, as summarized in Table 4. All adapters achieved fast inference times, with most completing predictions in under 0.1 seconds, reinforcing their suitability for real-time or resource-constrained clinical settings. SVM was the exception, with a longer but still acceptable inference time of 1.909 seconds. For training, KNN required negligible time (0.004 s), while LR (5.717 s) and MLP (5.147 s) provided a good balance between speed and performance. SVM (41.555 s) and RF (32.361 s) incurred higher training costs but less than one minute, remained within acceptable limits for efficient deployment.
@@ 4 Discussion
This study evaluated the diagnostic performance and computational efficiency of lightweight adapter classifiers trained on embeddings from pre-trained foundation models for multi-class classification of radiography images, with a focus on tube placement verification. Results show that adapter models effectively utilize these embeddings, achieving high mAUC values across multiple configurations. Notably, the combination of MedImageInsight embeddings with an SVM adapter achieved an mAUC of 93.85%, outperforming all other foundation model embeddings. In contrast, Med-Flamingo embeddings yielded substantially lower mAUC values, underscoring the importance of selecting embeddings that are well-aligned with the radiographic classification task.
Table 4: Average training and inference times (in seconds) on CPU for each adapter model
@@
An especially notable finding is the exceptional performance of MedImageInsight, despite it not being limited to radiography data. Unlike Rad-DINO and CXR-Foundation, which are tailored specifically to radiography and chest X-rays, MedImageInsight leverages data from a broad range of medical imaging modalities. This cross-modal training may enable the model to capture richer semantic features and shared medical imaging patterns, ultimately enhancing its representation power for radiographic tasks. These observations suggest that while task-specific domain adaptation is beneficial, diversity and scale across medical domains can provide additional gains, potentially offering more generalizable and semantically meaningful embeddings that transfer effectively across tasks.
@@
Despite achieving the best overall performance, the MedImageInsight + SVM combination struggled to distinguish between ETTube, NGTube, and VLine, as evidenced by the confusion matrix in Fig. 5. This issue is also apparent in the t-SNE visualization of MedImageInsight embeddings shown in Fig. 3, where these three classes are not clearly separated compared to others. Such overlap suggests that while the foundation model captures general semantic features, it may not preserve the fine-grained details needed to differentiate visually similar tube placements. These findings highlight a limitation of current multi-modal foundation models in retaining subtle cues critical for tasks involving highly similar anatomical structures or device orientations.
In the fairness analysis of MedImageInsight embeddings, the performance differences between genders across all adapters were minimal, with the largest observed gap being just 2.0%. Similarly, the variability across age groups was modest, with a maximum difference of 2.9%. However, a noticeable performance decrease was identified for the senior age group (80–100 years) within the Pneumothorax diagnostic class, particularly when using LG and MLP adapters, only 53.5% and 58.0% respectively shown in Table 3. This reduction may be attributable to the limited senior Pneumothorax patients in the test dataset, only three patients, suggesting more data are needed for accurate performance estimation.
@@
The results of our investigation provide compelling evidence that foundation model embeddings, especially those derived from domain-specific models such as MedImageInsight, can be harnessed to train efficient and accurate adapter classifiers. The observed superiority of the SVM adapter across most configurations further emphasizes the potential of combining pre-trained embeddings with lightweight, traditional machine learning classifiers[16, 17]. Moreover, the computational efficiency demonstrated—where training and inference times remained within seconds for most models and under one minute even for the more computationally demanding classifiers—supports the feasibility of integrating these methods into clinical workflows where rapid decision-making is essential.
@@
Despite these promising findings, several limitations must be acknowledged. Although the dataset encompassed a diverse range of clinical scenarios pertinent to tube placement, future studies should aim to validate these findings in prospective cohorts and across different clinical settings. Additionally, while our study focused mainly on tube placement verification—a task of significant clinical relevance—further research is warranted to assess the applicability of this methodology to other radiographic classification challenges.

### v2

@@ 3 Results
KNN F 88.9 89.5 97.2 91.3 88.4 97.3 90.3 91.8
M 87.1 86.0 96.1 88.6 90.8 96.1 86.7 90.2
Difference 1.8 3.5 1.2 2.7 -2.4 1.1 3.7 1.6
LG F 92.4 91.1 97.6 94.7 91.9 97.7 92.4 94.0
M 90.4 87.5 97.2 92.6 92.5 96.7 89.1 92.3
Difference 2.0 3.6 0.5 2.0 -0.6 1.0 3.3 1.7
SVM F 93.0 91.3 97.9 94.6 91.1 98.0 92.6 94.1
M 90.1 87.7 97.6 91.9 92.6 97.0 89.5 92.4
Difference 2.9 3.6 0.3 2.7 -1.5 1.0 3.1 1.7
RF F 91.0 90.4 97.4 94.4 90.9 97.4 91.5 93.3
M 89.3 86.9 96.8 92.3 92.0 96.6 88.3 91.8
Difference 1.7 3.5 0.6 2.1 -1.1 0.8 3.1 1.5
MLP F 92.7 91.5 97.2 95.3 92.7 98.1 91.8 94.2
M 90.4 87.5 97.7 91.7 92.3 97.0 89.8 92.3
Difference 2.3 4.0 -0.5 3.7 0.4 1.1 2.0 1.8
    KNN = k-nearest neighbors, LG = logistic regression, SVM = support vector machine, RF = random forest, MLP = multi-layer perceptron, ETTube=endotracheal tube, NGTube=nasogastric tube, RFrac=rib fracture, VLine=vascular lines, F=female patients (3641 in total), M=male patients (5195 in total).
Table 3: Age fairness assessment of AUC performance (%) from adapters trained by MedImageInsight embeddings. No significant difference was observed across age groups by a Kruskal-Wallis H-Test
@@
KNN (0, 20] 94.4 79.5 97.0 98.0 94.1 96.1 73.6 90.4
(20, 40] 86.7 89.6 96.4 89.1 89.5 95.8 89.6 91.0
(40, 60] 89.1 87.5 97.2 88.4 89.3 97.6 89.0 91.2
(60, 80] 86.7 85.7 96.3 89.8 90.9 96.6 88.7 90.6
(80, 100] 90.3 84.3 97.0 85.7 80.1 94.4 89.8 88.8
std 3.2 3.8 0.4 4.6 5.2 1.2 7.0 0.9
LG (0, 20] 93.3 84.1 98.4 98.8 97.0 99.0 81.2 93.1
(20, 40] 92.1 91.5 97.2 96.0 92.8 97.1 92.7 94.2
(40, 60] 91.6 88.9 97.7 92.3 90.5 97.6 90.5 92.7
(60, 80] 90.8 87.1 97.2 92.8 93.7 97.2 90.5 92.7
(80, 100] 93.3 87.9 97.3 86.2 80.9 96.3 93.2 90.7
std 1.1 2.7 0.5 4.7 6.1 1.0 4.9 1.3
SVM (0, 20] 97.9 84.5 98.4 95.9 97.1 99.1 82.0 93.6
(20, 40] 91.6 92.0 97.7 96.7 93.1 97.4 93.0 94.5
(40, 60] 91.2 88.8 97.9 90.8 89.6 97.9 90.5 92.4
(60, 80] 90.9 87.6 97.6 92.9 93.5 97.6 91.1 93.0
(80, 100] 92.7 88.6 97.7 86.6 80.9 96.1 93.4 90.9
std 2.9 2.7 0.3 4.1 6.2 1.1 4.6 1.4
RF (0, 20] 86.1 83.8 97.7 99.4 97.2 98.5 79.9 91.8
(20, 40] 90.5 90.9 97.2 95.5 92.1 97.0 91.8 93.6
(40, 60] 90.3 88.1 97.6 91.7 89.5 97.7 89.7 92.1
(60, 80] 90.3 86.5 96.8 92.8 92.6 97.3 90.0 92.3
(80, 100] 90.1 87.3 97.0 86.2 84.0 93.7 92.4 90.1
std 1.9 2.5 0.4 4.9 4.8 1.8 5.1 1.2
MLP (0, 20] 95.4 84.9 98.5 98.9 99.0 96.8 84.6 94.0
(20, 40] 90.7 92.9 97.8 97.7 91.3 97.8 93.5 94.5
(40, 60] 89.0 87.8 98.2 90.9 89.6 98.0 89.7 91.9
(60, 80] 92.2 87.7 96.2 92.3 95.4 97.2 90.4 93.1
(80, 100] 94.5 89.9 97.8 83.1 86.6 96.1 93.6 91.6
std 2.7 2.9 0.9 6.3 4.9 0.8 3.6 1.3
    ETTube = endotracheal tube, NGTube = nasogastric tube, RFrac = rib fracture, VLine = vascular lines, std = standard deviation. A total of 640 patients with age in the range of (0, 20], 909 patients with age in the range of (20, 40], 2161 patients with age in the range of (40, 60], 2525 patients with age in the range of (60, 80], and 418 patients with age in the range of (80, 100] are involved in the analysis. Patients older than 100 are unidentified and are excluded from analysis.
Beyond fairness, computational efficiency was assessed via CPU-based training and inference times, as summarized in Table 4. All models demonstrated low inference latency (<0.4 s), except SVM (2.764 s), which remained acceptable for real-time or resource-constrained clinical settings. KNN required negligible training time (0.004s), while MLP (6.374) offered a favorable balance of speed and performance. Although SVM (39.303s) and RF (155.322s) incurred higher training costs, all models remained within practical deployment limits.
@@ 4 Discussion
This study evaluated the diagnostic performance and computational efficiency of lightweight adapter classifiers trained on embeddings from pre-trained foundation models for multi-class classification of radiography images, with a focus on tube placement verification and findings. Results show that adapter models effectively utilize these embeddings, achieving high mAUC values across multiple configurations. Notably, the combination of MedImageInsight embeddings with an SVM or MLP adapter achieved an mAUC of 93.1%, outperforming all other foundation model embeddings. In contrast, Med-Flamingo embeddings yielded substantially lower mAUC values, underscoring the importance of selecting embeddings that are well-aligned with the radiographic classification task.
Table 4: Mean (standard deviation) of the training and inference time in seconds on CPU for each adapter model
@@
An especially notable finding is the exceptional performance of MedImageInsight and MedSigLIP, despite not being limited to radiography data. Unlike Rad-DINO and CXR-Foundation, which are tailored specifically to radiography and chest X-rays, MedImageInsight and MedSigLIP leverage data from a broad range of medical imaging modalities. This cross-modal training may enable the model to capture richer semantic features and shared medical imaging patterns, ultimately enhancing its representation power for radiographic tasks. These observations suggest that while task-specific domain adaptation is beneficial, diversity and scale across medical domains can provide additional gains, potentially offering more generalizable and semantically meaningful embeddings that transfer effectively across tasks. The Wilcoxon signed-rank test results further demonstrated that MedImageInsight embeddings were superior to other foundation models in this task.
@@
Despite achieving the best overall performance, the MedImageInsight + SVM combination struggled to distinguish between ETTube, NGTube, and VLine, as evidenced by the confusion matrix in Fig. 5. This issue is also apparent in the t-SNE visualization of MedImageInsight embeddings shown in Fig. 3, where these three classes are not clearly separated compared to others. Such overlap suggests that while the foundation model captures general semantic features, it may not preserve the fine-grained details needed to differentiate visually similar tube placements, which may be related to the pooling operation for the embedding extraction. These findings highlight a limitation of current multi-modal foundation models, particularly in generating embeddings from spatial features through simple pooling, which might obscure subtle cues that are critical for tasks involving highly similar anatomical structures or device orientations.
@@
The results of our investigation provide compelling evidence that foundation model embeddings, especially those derived from domain-specific models such as MedImageInsight, can be harnessed to train efficient and accurate adapter classifiers. The observed superiority of the SVM adapter across most configurations further emphasizes the potential of combining pre-trained embeddings with lightweight, traditional machine learning classifiers[16, 17]. Moreover, the computational efficiency demonstrated—where training and inference times remained within seconds for most models and around three minutes even for the more computationally demanding classifiers—supports the feasibility of integrating these methods into clinical workflows where rapid decision-making is essential.
@@
Despite these promising findings, several limitations should be acknowledged. Although the dataset encompassed a diverse range of clinical scenarios pertinent to tube placement, future studies should aim to validate these findings in prospective cohorts and across different clinical settings. Additionally, many chest X-ray findings commonly co-occur, and a multi-label classification framework would more closely reflect clinical reality. The current single-label multi-class approach is to simplify the problem formulation and to establish a clear, reproducible benchmark task suitable for a wide range of modeling techniques. Therefore, this work is intended to serve as an initial step toward more clinically representative, multi-label classification tasks, which will be explored in future studies. Furthermore, while our study focused mainly on tube placement verification and related findings—a task of significant clinical relevance—further research is warranted to assess the applicability of this methodology to other radiographic classification challenges.